Upper Division of Lower Company acquired an asset with a cost of $590,000 and a four-year life. The cash flows from the asset, considering the effects of inflation, were scheduled as follows.

Year Cash Flow
1 $ 210,000
2 265,000
3 285,000
4 325,000

The cost of the asset is expected to increase at a rate of 20 percent per year, compounded each year. Performance measures are based on beginning-of-year gross book values for the investment base. Ignore taxes.

Required:

a. What is the ROI for each year of the asset's life, using a historical cost approach?

b. What is the ROI for each year of the asset's life if both the investment base and depreciation are determined by the current cost of the asset at the start of each year?
Answer #1

Answer:

a.

Year    ROI   
1 10.59%
2 19.92%
3 23.31%
4 30.08%

Working :

Cost of the asset 590000
Life of the Asset 4 years
Annual Depreciation 147500 =590000/4
Year Cash Flow Depreciation Net Income
(Cash Inflow - Depreciation)
ROI
(Net Income / Cost of Asset )
1 $ 210,000.00 $ 147,500.00 $   62,500.00 10.59% =62500/590000
2 $ 265,000.00 $ 147,500.00 $ 117,500.00 19.92% =117500/590000
3 $ 285,000.00 $ 147,500.00 $ 137,500.00 23.31% =137500/590000
4 $ 325,000.00 $ 147,500.00 $ 177,500.00 30.08% =177500/590000

b.

Year ROI
1 10.59%
2 3.61%
3 14.05%
4 18.56%

Working :

Year Asset Cost Cash Flow Depreciation Net Income
(Cash Inflow - Depreciation)
ROI
(Net Income / Cost of Asset )
1 $ 590,000.00 $ 210,000.00 $ 147,500.00 $   62,500.00 10.59% =62500/590000
2 $ 1,239,000 $ 265,000.00 $ 309,750.00 $ 44,750.00 3.61% =44750/1239000
3 $ 2,601,900 $ 285,000.00 $ 650,475.00 $ 365,475.00 14.05% =365475/2601900
4 $ 5,463,990 $ 325,000.00 $ 1,365,997.5 $ 1,013,997.5 18.56% =1013997/5463990